Director's Time

Observers: B.Rothberg, O. Kuhn, M. Wagner (Tucson)

ISA: O.Kuhn (Tucson)
Telescope Operator: S.Allanson (LBTO)
AO Support: J.Christou
Telescope Support: J.Little (LBTO)
Software Support: C.Biddick
Instrument Support: J.Power (LBTO)
Engineering Support: M.Gusick

Summary

MODS2 Commissioning with some Science Time

Details

UT 05:50 - MODS1 observing FF1456+3337 -Rothberg science (mods1b.20150512.0002-0005, mods1r.20150512.0006-0013)

UT 06:04 - avg seeing guider 0.89"

UT 06:16 - avg seeing guider 0.89" DIMM 1.03" - probably thin cirrus, clouds on NE sky

UT 06:55 - exposure 3 during a zenith transit - no sign of continuum - lost object - unsure if -10" dither or transit caused this to happen. Waiting for transit to complete before resending acquisition (mods1b.20150512.0004,mod1r.20150512.0008)

UT 07:03 - Resending acquisition

UT 07:11 - still needed to send a -0.246" (2pix) offset

UT 07:14 - starting science

UT 07:38 - Did 1 exposure - moving to Feige 67

UT 07:43 - have to do an unwrap - other specstd is transiting so would cause a problem to observe now

UT 07:47 - at Feige 67 airmass 1.392

UT 07:54 - star didn't end up in slit - not sure why - resending acquisition - correction - reason was acq image was 3s which is too short for a well-lit thru slit image. This caused the center of the slit finding algorithm to think the X coord was 456 not 501.

UT 08:03 - Feige 67 integration start airmass 1.483-1.564 (mods1b.20150512.0006-0007,mods1r.20150512.0021-0022)

UT 08:17 - MODS1 done for now.

UT 09:30 - going back to science - pseudo-monocular - MODS1 + LBC-R

UT 09:31 - PDS 456 - PI Kuhn

UT 09:46 acqMODS PDS456k23.acq

UT 09:54 on-target, dofpia, /first, /redonly (seeing on guider is 1.1-1.4" and some cirrus).

UT 09:57 modsAlign offsetxy -3.562 12.552 rel to put quasar at Y~640.

UT 10:00 execMODS PDs456_900sec.obs (will update pointing, go at 0 0 and then offset 0 .15, offset -0.15, taking 15-min exp at each pos). Take a 4th exposure at offset 0 0.30 abs (via gui).

UT 10:00 dofpia needs attention, as initial pupils (T=3.8 C) look practically in-focus. Needed a manual offset of -10,000. From there dofpia, /first, /redonly iterated.

UT 10:10 ./loop_observe.py -p /home/lbto/kuhn/PDS456/LBC/ -o 'PDS456riz.ob PDS456riz_longer.ob' All in-focus images are out of focus! And red LBC Sci is ahead by 2 minutes.

UT 10:26 No instrument offsets are being sent, although active optics are being changed as dofpia runs. Barry doing an lbckill/lbcstart. This did not help.

UT 10:41 We conclude that LBC is sending active optics offsets to PSF, but it is not sending instrument offsets to PSF.

UT 11:09 MODS1 exposure ended.

UT 11:09 Steve restarted PSF Right. Preset is still ongoing. This did not help - active optics being sent to PSFR, but no instrument offsets. Barry rebooting (for the first time) the CMU. This did the trick. Wrote 5564, cross-ref 2214.

UT 11:20 dofpia converged. A few exposures on PDS456. Superfoc RB_rV_sf7.ob in twilight.

UT 12:04 - MODS-1 calibrations

UT 12:05 - slitless pixel flats (0012-0021 blue, 0030-0034 red)

UT 12:22 - Arc lamps (0022-0024 blue, 0035-0037 red)

UT 12:31 - Slt flats (0.6" and 1.0") (0025-0030,0031-0036 blue, 0038-0043,0044-0049 red)

UT 13:10 - 8k Biases (0037-0041 blue, 0050-0054 red)

-- %USERSIG{BarryRothberg - 2015-05-12}%

Comments

 
Topic revision: r4 - 12 May 2015, BarryRothberg
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback